### v1

@@ Abstract
In this work we present a novel multi-view photometric stereo (PS) method. Like many works in 3D reconstruction we are leveraging neural shape representations and learnt renderers. However, our work differs from the state-of-the-art multi-view PS methods such as PS-NeRF [40] or SuperNormal [1] we explicity leverage per-pixel intensity renderings rather than relying mainly on estimated normals.
We model point light attenuation and explicitly raytrace cast shadows in order to best approximate each points incoming radiance. This is used as input to a fully neural material renderer that uses minimal prior assumptions and it is jointly optimised with the surface. Finally, estimated normal and segmentation maps can also incorporated in order to maximise the surface accuracy.
Our method is among the first to outperform the classical approach of DiLiGenT-MV and achieves average 0.2mm Chamfer distance for objects imaged at approx 1.5m distance away with approximate $400\times 400$ resolution. Moreover, we show robustness to poor normals in low light count scenario, achieving 0.27mm Chamfer distance when pixel rendering is used instead of estimated normals.
Keywords:
Multiview Photometric Stereo Neural Rendering Neural Point Light Rendering
@@ 1 Introduction
Photometric Stereo (PS) is a long standing and very important problem in the field of Computer Vision problems. While early PS works [39, 11, 27, 13] primarily tackled the estimation of normals from a single view image, their potential was really unlocked by binocular [18, 4, 38, 24] and multi-view [5, 46, 34, 28, 21, 40] stereo setups as it allowed for a more accurate recovery of shape and not only normals, opening multiple applications such as general 3D reconstruction of objects for novel-view rendering, relighting or material editing [40], robot interaction, quality control in manufacturing or industrial conveyor belt scanning.
@@
Figure 1: In this figure we demonstrate the fragility of relying mainly on estimated normals (using [9]) for deep learning based multi-view photometric stereo. The first column shows the ground truth of Buddha object from DiLiGenT-MV benchmark. The second column shows the estimated normals using only 6 out of 96 lights available (view 1 mean average normal error is $13.5^{\circ}$). Using such normals leads to a large error in reconstruction (0.49mm). If intensities are used instead an error of 0.29mm is achieved. The right column shows the result using all lights and we note that we outperform DiLiGenT-MV [21], PS-NeRF [40] and MVAS [2] and only loose to SuperNormal [1], as shown in Table 2.
@@
Along with increasing the number of views Photometric Stereo undergone another important change by moving from classical non-linear optimisation enabled inverse graphics approaches (for single view [10], binocular [18], multi-view [28, 21]) to neural network (e.g. [11, 26]) and in particular neural representation enabled inverse graphics approaches (for single view [8], binocular [24] and multi-view [40, 1]). Despite the latter methods (especially [1]) achieving impressive accuracy on DiLiGenT-MV [21] benchmark their approaches are somewhat unsatisfactory. In particular,  [1] does not explicitly use image intensities to optimise for shape and it is fully reliant on per-view normal maps. Whereas, PS-NeRF [40] only uses average intensity during the surface optimisation stage and thus leaves most of the photometric information unused.
@@
In this work we provide the first neural multi-view photometric stereo approach which fully leverages the availability of pixel intensity information for estimating 3D shape from Photometric Stereo images (see Figure 5). We achieve this by explicitly modeling the incident light from point light sources to leverage intensity based shape optimisation over purely normal driven shape optimisation [40, 1] which is fragile to incorrectly estimated normals especially in cases of very few available lights.
We model point light attenuation, explicitly raytrace cast shadows and also optimise a fully neural material renderer. Thus we are able to achieve competitive reconstruction accuracy using only 6 lights and match the SOTA performance when all lights and normal map information is also fused.
@@
The remainder of this paper is organised as follows. Section 2 discusses the related works in Photometric Stereo and Multi-View Stereo. It is followed by a description of our method in Section 3. The experimental setup and experiment results are described in Sections 4 and 5 respectively.
@@ 2 Related Work
Single view photometric stereo. The first successful deep learning based single view PS was CNN-PS [11] which was extended by [26] and [27] to be applicable to general calibrated point like configurations. Other works like [7, 5, 14] have used material reflectance priors (using specific BRDFs like Lambertian or Ward) for single view normal prediction. Other recent approaches have also tackled a more uncalibrated setting like [3, 20, 19, 41] and more recently the idea of universal uncalibrated PS [12, 13] and [9].
Multi-view photometric stereo. Classical optimisation approaches have used triangle meshes [34] or sign distance function based parameterisations [28, 33, 47] to tackle the multi-view PS problem, under diffuse reflectance. More advanced methods such as [21] are more applicable to more general materials.
Neural surfaces. Recently, neural surface approaches have become very popular from the introduction of NeRF [32] and its extension to neural SDF parameterisations in [43, 42]. The first approaches to used neural SDFs for specifically the multi-view PS problem are [17, 15, 16, 45], PS-NeRF [40] and SuperNormal [1]. However, these approaches are mainly relying on normal fusion and do not attempt to update the neural surface though rendering. Other recent approaches have advanced the sophistication of the rendering methods to be more structured and thus respect the physics of light reflection more, like Ref-NeRF [37], Neuralangelo [22], NERO [23] and NeILF++ [44] but non of these methods has yet to be applied to PS problem, especially lacking the ability to model point light illumination. Finally it is worth mentioning [8] who introduced the idea of a infinitely differentiable surface (SIREN [36]) with Lambertian rendering for the single view PS and [25], which extended this method to the binocular setting and also added a fully learnable general material renderer.
We borrow the material renderer from [25] while extending the approach to work in the multi-view setting using an SDF paremeterisation similar to [43, 42].
@@ 3 Method
This section describes our method for solving the point light multi-view photometer stereo. A high level overview is also shown in Figure 2. Our method is primarily an inverse neural rendering method. Section 3.1 describes the assumed irradiance model. Section 3.2 describes the underlying neural surface parametersiation and Section 3.4 describes training losses used.
Figure 2: High level schematic of our overall method. Single view PS is used to obtain normal maps. Training the SDF with normal loss obtains a rough surface which is then refined with full volumetric rendering. The rendering procedure is explored in Figure 3. The second row also shows the GT and render images (as grayscale), the rendering error (with red =0.1) as well as the computed shadow map.
@@ 3.1 Irradiance equation
We now explain the assumed irradiance of a world point $\boldsymbol{\mathbf{X}}$ with surface normal $\boldsymbol{\mathbf{N}}$ and albedo $\rho$. We assume point light sources $m$ at positions $\mathbf{P}_{m}$ which generate variable lighting vectors $\mathbf{L}_{m}=\mathbf{P}_{m}-\mathbf{X}$. In addition, point light propagation results to the following attenuation factor $a_{m}=\frac{\phi_{m}}{||\mathbf{L}_{m})||^{2}}$ where $\phi_{m}$ is the intrinsic brightness of the light source. We note that the literature [31] usually also assumes angular dissipation factor but these calibration numbers are unavailable for DiLiGenT-MV [21] therefore we opt for the simpler, perfect point light source model. Thus, the reflected intensity of that point (not include $\boldsymbol{\mathbf{X}}$ for clarity) for the $m$ th light source $i_{m}$ is:
$i_{m}=s_{m}a_{m}\rho B(\boldsymbol{\mathbf{N}},\boldsymbol{\mathbf{L}}_{m},\boldsymbol{\mathbf{V}})+i_{r,m}$ (1)
B is assumed to be a general BRDF, $s\in\{0,1\}$ is an indicator variable to account for cast shadows that completely block direct reflectance. $i_{r,m}$ corresponds to indirect reflectance (e.g. self reflections) and we will further make the simplifying assumption that it is also proportional to the surface albedo as well as the light source attenuation. Finally, we note that self reflection occur on concave parts of an objects surface which also correlate with more cast shadows therefore we make the following assumption $i_{r}\propto(a\rho\sum(1-s))$. Indeed, the term ambient occlusion is usually used to describe the total obliqness of a surface point and this is approximated with the number of shadows. To simplify the learning problem, we assume that $i_{r}$ is constant between different lights thus $i_{r,m}=a\rho\sum sR(\boldsymbol{\mathbf{X}})$. This assumption is further justified empirically on the supplementary by empirical measurements on Blender-rendered data.
@@ 3.2 Neural SDF
Figure 3: Visualisation of our volume rendering approach. Two rays with multiple ray samples $\boldsymbol{\mathbf{X}}_{ri}$, and $\boldsymbol{\mathbf{Y}}_{ri}$ and shown. The average intersection points $\boldsymbol{\mathbf{X}}_{A}$, and $\boldsymbol{\mathbf{Y}}_{B}$ are also shown as these are used to ray trace cast shadows (towards the light source at position $\boldsymbol{\mathbf{P}}$ with brightness $\phi$). Cast shadow samples are marked as $\boldsymbol{\mathbf{X}}_{si}$, and $\boldsymbol{\mathbf{Y}}_{si}$ respectively. Note that points that significantly contribute to the total rendering (though the accumulated opacity) are coloured blue and points that do not (because they are outside of the surface or occluded) are marked red. For these red points, the free space regulariser is used. For shadow sample points rendering is not performed and so are marked black. Note that the average intersection points are only used to guide shadows so they are not rendered either. Finally, for the $\boldsymbol{\mathbf{X}}_{r2}$ ray sample point, normal $\boldsymbol{\mathbf{N}}$, lighting $\boldsymbol{\mathbf{L}}$ and viewing vectors $\boldsymbol{\mathbf{V}}$ (that are used for rendering) are showing with respective colors of red,green and blue.
@@
Geometry parameterisation. Following the work of other neural volumetric approaches [43, 42], we parameterise the scene geometry as the zeroth level set of an implicit function $F$ corresponding to the Signed Distance Field, $d=F(\boldsymbol{\mathbf{X}})$. The unknown function $F$ is a deep neural network and the objective is to optimise its weights. We note that the SDF of any arbitrary geometry is always continuous, almost everywhere differentiable and satisfies the Eikonal equation of unit magnitude gradient $||\nabla F(\boldsymbol{\mathbf{X}})||=1$. Finally, we note that for surface points (where $F(\boldsymbol{\mathbf{X}})=0$), the surface normal is the gradient of the SDF, i.e. $\boldsymbol{\mathbf{N}}(\boldsymbol{\mathbf{X}})=\nabla F$. This allows for a direct SDF to rendering relationship and thus allowing to train the SDF through rendering loss from the initial photometric stereo images. In addition, if surface normal maps are available (e.g. from single view estimation networks) they can also be used as an addition training signal. We use the SIREN architecture [36] which is a MLP with sinusoidal activation functions and that guarantees that the surface is infinitely differentiable thus can be easily recovered from its derivatives. Indeed, [8, 25, 44] have used SIREN networks for single and binocular view PS as well as multi-view stereo so we are a direct extension to them specifically for the MV-PS under point light sources.
@@
Ray sampling. We follow a volumetric sampling and rendering method similar to VolSDF [42] where the neural SDF is queried in multiple samples on outgoing rays from each image foreground pixel. For each pixel an estimate of the depth is available (initialised from single view PS and occasionally updated during training) and thus most of the samples are concentrated around that depth. However, to allow the surface to evolve and to minimse free space artefacts, additional samples are also sampled in a bigger depth range. As each point, the Laplace density function (see VOLSDF [42] ) is used to convert from SDF values to transparency and alpha blending is used to accumulate depth, normals and rendered intensity at each ray using the standard approach.
@@
Thus, transparecny $\alpha=\text{exp}(-t\delta r)$ with $\delta r$ being the distance along the ray and so the intersection $\boldsymbol{\mathbf{X}}_{A}$ is computed as a weighted sum of ray samples $\boldsymbol{\mathbf{p}}=\sum_{i}(w_{i}\boldsymbol{\mathbf{r}}_{i})$, with the sample weights $w_{i}$ corresponding to the accumulated opacity. We note that this average points $\boldsymbol{\mathbf{X}}_{A}$ is further used to compute cast shadows but it is not directly render. Similar averaging is used for the rendered intensity as well as surface normal which are used to compute losses. The ray sampling process is shown in Figure 3.
@@
Learned BRDF. We follow the approach of [25] where the BRDF is also parameterised as another SIREN network and thus completely learned from the data. This assumes that the material properties are uniform around the target scene except for a scalar albedo variation. We emphasise that we chose to perform grayscale intensity rendering instead of full RGB as this is expected to minimise the synthetic to real gap. Indeed, real RGB images are usually acquired with demosaicing of single intensity values and this procedure is usually optimised to best recover intensity not colour (e.g see [6] used by OpenCV).
@@
To minimise over-fitting, the material BRDF network receives as input only the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{V}}$. In addition, to simplify the learning problem, we following the principles described in the MERL real material database [30]. For that, the half vector $\boldsymbol{\mathbf{H}}=\frac{\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}}{|\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}|}$ is first computed and the input to the network is the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{H}}$. Finally, we note that the final activation of SIREN part of the BRDF network is exponential and there is a post multiplication with an $\boldsymbol{\mathbf{N}}\cdot\boldsymbol{\mathbf{L}}$ factor such as the networks learns a multiplicative factor over the diffuse reflectance, so thus we parameterise:
@@
Shadow estimation. Handling cast shadows is imperative for any successful photometric stereo method. To estimate cast shadows for a point, we raytrace from that point to the light source following the direction of the lighting vectors $\boldsymbol{\mathbf{l_{m}}}$ computed above. For each ray we take 16 random samples $j$ $\in[2\text{mm},50\text{mm}]$. For all these points, we query the SDF network and accumulate opacity following the same volumetric rendering procedure, i.e $s=\prod_{j}(1-\alpha_{j})$. We note that this shadow computation procedure is very computationally expensive therefore it cannot be computed for all the rendered points. Instead, we only compute it for the average intersection point of each ray ($\boldsymbol{\mathbf{X}}_{A}$ is Figure 3) and assume it constant for all the other ray samples.
Indirect reflectance. Finally, the indirect reflectance $i_{r,m}$ (Equation 1) is approximated as $i_{r,m}=a\rho\tilde{s}R(\boldsymbol{\mathbf{X}})$. $\tilde{s}$ is compute as the mean value of one minus shadows (over the number of lights) and its a measure of the local surface concavity; $\tilde{s}=0$ corresponds to no shadows and hence completely convex local geometry therefore no intereflection, $\tilde{s}=1$ corresponds to all lights being shaded and hence maximum concavity therefore we allow for maximum learned intereflection. Finally $R(\boldsymbol{\mathbf{X}})$ is queried from another SIREN network and we note that it is parameterised to be constant on all lights.
Some visualisation of the generated rendering for synthetic and real expeiments are shown in Figure 4.

### v2

@@ Abstract
In this work we present a novel multi-view photometric stereo (MVPS) method. Like many works in 3D reconstruction we are leveraging neural shape representations and learnt renderers. However, our work differs from the state-of-the-art multi-view PS methods such as PS-NeRF [44] or Supernormal [4] in that we explicitly leverage per-pixel intensity renderings rather than relying mainly on estimated normals.
We model point light attenuation and explicitly raytrace cast shadows in order to best approximate the incoming radiance for each point. The estimated incoming radiance is used as input to a fully neural material renderer that uses minimal prior assumptions and it is jointly optimised with the surface. Estimated normals and segmentation maps are also incorporated in order to maximise the surface accuracy.
Our method is among the first (along with Supernormal [4]) to outperform the classical MVPS approach proposed by the DiLiGenT-MV benchmark and achieves average 0.2mm Chamfer distance for objects imaged at approx 1.5m distance away with approximate $400\times 400$ resolution. Moreover, our method shows high robustness to the sparse MVPS setup (6 views, 6 lights) greatly outperforming the SOTA competitor (0.38mm vs 0.61mm), illustrating the importance of neural rendering in multi-view photometric stereo.
@@ 1 Introduction
Photometric Stereo (PS) is a long standing and very important problem in the field of Computer Vision. While early PS works [43, 14, 31, 16] primarily tackled the estimation of normals from a single view image, the value of PS was unlocked by binocular [21, 7, 42, 28] and multi-view [8, 50, 38, 32, 24, 44] stereo setups as it allowed for accurate recovery of shape and not only normals. This, in turn, opened many applications such as general 3D reconstruction, novel-view rendering, relighting and material editing [44], as well as robot interaction, quality control in manufacturing and industrial conveyor belt scanning.
@@
Figure 1: In this figure we demonstrate the fragility of relying mainly on estimated normals (using [12]) for deep learning based sparse multi-view photometric stereo when 6 views out of 20 and 6 light sources out of 96 available are used. The first column shows the ground truth of Buddha object from dense MVPS DiLiGenT-MV benchmark. The following two columns show the estimated normals and corresponding error map using only 6 out of 96 lights available (view 1 mean average normal error is $9.9^{\circ}$, saturated red color corresponds to $5^{\circ}$ error). Using such normals leads to a large error in reconstruction error using our method when pixel intensities are not leveraged (0.51mm) and previous SOTA dense MVPS method Supernormal [4] (0.67mm). If pixel intensities are used along with estimated normals (column 6) a significantly smaller error of 0.35mm is achieved. The final two columns show the error maps of estimated shapes when all available views and lights are used where Supernormal [4] achieves a similar reconstruction error as our method (0.21mm vs 0.19mm). Similar dynamics apply to other DiLiGenT objects as shown in Tables 2 and 3, strongly motivating for explicit pixel intensity modeling in MVPS methods. Note here the errors are computed as Chamfer distance while the visualisation only shows errors from reconstruction to ground truth mesh for each ground truth mesh surface point. Note dark red corresponds to 1mm error in the shape error illustrations (columns 3-8).
@@
Along with increasing the number of views Photometric Stereo undergone another important change by moving from classical non-linear optimisation enabled inverse graphics approaches (for single view [13], binocular [21], multi-view [32, 24]) to neural network (e.g. [14, 30]) and in particular neural shape representation enabled inverse graphics approaches (for single view [11], binocular [28] and multi-view [44, 4]). However, despite the latter methods, especially [4], achieving impressive accuracy on DiLiGenT-MV [24] benchmark their approach to MVPS is somewhat incomplete as they do not attempt to directly explain (and learn to match) observed pixel-wise intensities. In particular,  [4] does not explicitly use image intensities to optimise for shape and it is fully reliant on per-view normal maps. Whereas, PS-NeRF [44] only uses average intensity during the surface optimisation stage and thus leaves most of the photometric information unused. It is important to note that Brahimi et. al. [1] , attempts to re-render the images however does not model cast shadows and uses the simplified Dinsey BRDF [3] which may not model all materials accurately.
@@
In this work we provide the first neural multi-view photometric stereo approach which fully leverages the availability of pixel intensity information for estimating 3D shape from Photometric Stereo images (see Figure 5). We achieve this by explicitly modeling the incident light from point light sources to leverage intensity based shape optimisation over purely normal driven shape optimisation [44, 4] which is fragile to incorrectly estimated normals especially in cases of very few available lights as shown Figure 1 and Section 5.
In more detail, we model point light attenuation and explicitly raytrace cast shadows in order to best approximate the incoming radiance for each point. The estimated incoming radiance is used as input to a fully neural material renderer that uses minimal prior assumptions and it is jointly optimised with the surface. Estimated normals and segmentation maps are also incorporated in order to maximise the surface accuracy. This allows us to achieve SOTA reconstruction accuracy (0.2mm) on original dense (20 views, 96 lights) DiLiGenT-MV [24] benchmark and signifcantly outperform (0.38mm vs 0.61mm) the best MVPS method [4] in a sparse setup of 6 views and 6 lights.
@@
The remainder of this paper is organised as follows. Section 2 discusses the related work in Photometric Stereo and Multi-View Stereo. It is followed by a description of our method in Section 3. The experimental setup and experiment results are described in Sections 4 and 5 respectively.
@@ 2 Related Work
Single view photometric stereo. The first successful deep learning based single view PS was CNN-PS [14] which was extended by [30] and [31] to be applicable to general calibrated point like configurations. Other works like [10, 8, 17] have used material reflectance priors (using specific BRDFs like Lambertian or Ward) for single view normal prediction. Other recent approaches have leveraged the power of recent transformer models and big synthetic datasets (often of tens of thousand of images) to tackle a weakly uncalibrated setting like [6, 23, 22, 45] and more recently fully uncalibrated single-view PS [15, 16] and [12]. However, despite the success of these methods in producing single view normal maps (as well as material maps), accurate shape reconstruction is still challenging.
Multi-view photometric stereo. To overcome the ill-posedness of single view photometric surface reconstruction, multi-view photometric stereo (MVPS) methods have leveraged information from multiple view and multiple lights. Classical optimisation approaches have used triangle meshes [38] or sign distance function based parameterisations [32, 37, 51] to tackle the multi-view PS problem, under diffuse reflectance. Methods, e.g. [24], were also developed for more general materials as well.
Neural surfaces. Recently, neural surface approaches have became very popular in tacking the 3D reconstruction problem. Early approaches include NeRF [36] and its first extensions to neural SDF parameterisations [47, 46]. The first methods which used neural SDFs specifically for the multi-view PS problem include [20, 18, 19, 49]. However, contrary to the direction of the neural inverse rendering literature, these approaches do not attempt to re-render the original photometric stereo images but rather some 2D derivates of the images such as normals or albedo maps. For example, Supernormal [4] only renders normal maps but achieves very fast training speed though patch parameterisation, as well as the use of the NERFACC [25] framework. PS-NeRF [44] renders normal and average intensity maps whereas RNb-Neus [2], uses normal and albedo maps to render virtual light images. Thus, all these methods are significantly reliant on single view PS networks and have no way to circumvent noisy estimates that are likely to happen in case of sparse lights and number of views as demonstrated in Figure 1.
Other recent neural rendering approaches have advanced the sophistication of the rendering methods to be more structured and thus respect the physics of light reflection more, like Ref-NeRF [41], Neuralangelo [26], NERO [27] and NeILF++ [48] but none of these methods has yet to be applied to PS problem, especially lacking the ability to model point light illumination. Finally it is worth mentioning [11] who introduced the idea of a infinitely differentiable surface (SIREN [40]) with Lambertian rendering for the single view PS and [29], which extended this method to the binocular setting and also added a fully learnable general material renderer.
We borrow the material renderer from [29] while extending the approach to work in the multi-view setting using an SDF parametrisation similar to [47, 46]. It is worth noting that Brahimi et. al, [1], also tackles the MVPS problem though physics-based per point rendering (using the Disney BRDF [3]). This approach is the most similar to us with major differences being that we use a fully neural material model, we explicitly ray-trace cast shadows and also employ supervision signal from single view photometric stereo normals. Thus we are able to outperform them with 0.25mm vs 0.34mm error (see Table 2).
@@ 3 Method
Figure 2: Schematic of our overall method. Single view PS is used to obtain normal maps. Training the SDF with normal loss obtains a rough surface which is then refined with full volumetric rendering. The rendering procedure is explained in Figure 3. The second row also shows the GT and render images (as grayscale), the rendering error (with red =0.1) as well as the computed shadow map.
This section describes our method for solving the point light multi-view photometer stereo. A high level overview is also shown in Figure 2. Our method is primarily an inverse neural rendering method. Section 3.1 describes the assumed irradiance model. Sections 3.2 and 3.3 describe the underlying neural surface parametrisation and its initialisation, respectively. Section 3.4 describes training losses used.
@@ 3.1 Irradiance equation
We now explain the assumed irradiance equation of a world point $\boldsymbol{\mathbf{X}}$ with surface normal $\boldsymbol{\mathbf{N}}$ and albedo $\rho$. We assume point light sources $m$ at positions $\mathbf{P}_{m}$ which generate variable lighting vectors $\mathbf{L}_{m}=\mathbf{P}_{m}-\mathbf{X}$. In addition, point light propagation results to the following attenuation factor $a_{m}=\frac{\phi_{m}}{||\mathbf{L}_{m}||^{2}}$ where $\phi_{m}$ is the intrinsic brightness of the light source. We note that the literature [35] usually also assumes angular dissipation factor but these calibration numbers are unavailable for DiLiGenT-MV [24] therefore we opt for a simpler, perfect point light source model. Thus, the reflected intensity of the point $\boldsymbol{\mathbf{X}}$ for the $m$-th light source $i_{m}$ is modelled as:
$i_{m}=s_{m}a_{m}\rho B(\boldsymbol{\mathbf{N}},\boldsymbol{\mathbf{L}}_{m},\boldsymbol{\mathbf{V}})$ (1)
Note, here $B(.)$ is assumed to be a general BRDF, $s_{m}\in\{0,1\}$ is an indicator variable to account for cast shadows that completely block direct reflectance. Also note, we assume that indirect reflectance (i.e. self reflections and ambient light) are negligible and can be ignored.
@@ 3.2 Neural SDF
Figure 3: Visualisation of our volume rendering approach. Two rays with multiple ray samples $\boldsymbol{\mathbf{X}}_{ri}$, and $\boldsymbol{\mathbf{Y}}_{ri}$ are shown. The surface-ray intersection points $\boldsymbol{\mathbf{X}}_{I}$, and $\boldsymbol{\mathbf{Y}}_{I}$ are also shown as they are used to ray trace cast shadows (towards the light source at position $\boldsymbol{\mathbf{P}}$ with brightness $\phi$). Cast shadow samples are marked as $\boldsymbol{\mathbf{X}}_{si}$, and $\boldsymbol{\mathbf{Y}}_{si}$ respectively. Note that points that significantly contribute to the total rendering (though the accumulated opacity) are coloured blue and points that do not (because they are outside of the surface or occluded) are marked red. For shadow sample points rendering is not performed and so are marked black. Note that the intersection points ($\boldsymbol{\mathbf{X}}_{I}$, and $\boldsymbol{\mathbf{Y}}_{I}$) are only used to guide shadows so they are not rendered either. Finally, for the $\boldsymbol{\mathbf{X}}_{r2}$ ray sample point, normal $\boldsymbol{\mathbf{N}}$, lighting $\boldsymbol{\mathbf{L}}$ and viewing vectors $\boldsymbol{\mathbf{V}}$ (that are used for rendering) are shown with respective colors of red,green and blue.
@@
Geometry parameterisation. Following the work of other neural volumetric approaches [47, 46], we parameterise the scene geometry as the zeroth level set of an implicit function $F$ corresponding to the Signed Distance Field, $d=F(\boldsymbol{\mathbf{X}})$, parameterised by a deep neural network. We note that the SDF of any arbitrary geometry is always continuous, almost everywhere differentiable and satisfies the Eikonal equation of unit magnitude gradient $||\nabla F(\boldsymbol{\mathbf{X}})||=1$. Finally, we note that for surface points (where $F(\boldsymbol{\mathbf{X}})=0$), the surface normal is the gradient of the SDF, i.e. $\boldsymbol{\mathbf{N}}(\boldsymbol{\mathbf{X}})=\nabla F$. This allows to train the SDF through rendering loss from the initial photometric stereo images. In addition, if surface normal maps are available (e.g. from single view estimation networks) they can also be used as an addition training signal. We use the SIREN architecture [40] which is a MLP with sinusoidal activation functions and that guarantees that the surface is infinitely differentiable thus can be easily recovered from its derivatives.
@@
Ray sampling. We follow a volumetric sampling and rendering method similar to VolSDF [46] where the neural SDF is queried in multiple samples on outgoing rays from each image foreground pixel. For each pixel an estimate of the depth is available (initialised from single view PS and occasionally updated during training) and thus most of the samples are concentrated around that depth. However, to allow the surface to evolve and to minimse free space artefacts, additional samples are also sampled in a wider depth range. Following VOLSDF [46], we use the Laplace density function to convert from SDF values $d$ to transparency $t$ as follows:
$t(d)=\big{(}0.5+0.5\cdot sign(d)(\exp(-|\beta d|)-1)\big{)}/\beta$ (2)
with $\beta$ being a trainable scalar constant controlling the sharpness of the distribution.
@@
Finally, alpha blending is used to accumulate depth, normals and rendered intensity at each ray using the standard approach, i.e. transparecny $\alpha=\text{exp}(-t\delta r)$ with $\delta r$ being the distance along the ray and so the surface-ray intersection point $\boldsymbol{\mathbf{X}}_{I}$ is computed as a weighted sum of ray samples $\boldsymbol{\mathbf{X}}_{I}=\sum_{i}(w_{i}\boldsymbol{\mathbf{X_{r}}}_{i})$, with the sample weights $w_{i}$ corresponding to the accumulated opacity. Note that intersection points $\boldsymbol{\mathbf{X}}_{I}$ are further used to compute cast shadows but are not directly rendered. Similar averaging is used to obtain the rendered intensity ($i_{I}=\sum_{i}(w_{i}i_{i})$) and surface normal ($\boldsymbol{\mathbf{N}}_{I}=\sum_{i}(w_{i}\boldsymbol{\mathbf{N_{r}}}_{i})$) both of which are used to compute losses. The ray sampling process is further explained in Figure 3.
@@
Learned BRDF. We follow the approach of [29] where the BRDF is also parameterised as another SIREN network and thus completely learned from the data. This assumes that the material properties are uniform around the target scene except for a scalar albedo variation. We emphasise that we chose to perform grayscale intensity rendering instead of full RGB rendering as this is expected to minimise the synthetic to real gap. Indeed, real RGB images are usually acquired with demosaicing of single intensity values and this procedure is usually optimised to best recover intensity not colour (e.g see [9] used by OpenCV).
@@
To minimise over-fitting, the material BRDF network receives as input only the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{V}}$. In addition, to simplify the learning problem, we follow the principles described in the MERL real material database [34]. To achieve this, the half vector $\boldsymbol{\mathbf{H}}=\frac{\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}}{|\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}|}$ is first computed and the input to the network is the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{H}}$. Finally, we note that the final activation of SIREN part of the BRDF network is exponential and there is a post multiplication with an $\boldsymbol{\mathbf{N}}\cdot\boldsymbol{\mathbf{L}}$ factor so that the BRDF network learns a multiplicative factor over the diffuse reflectance, parameterised as follows:
@@
Shadow estimation. Handling cast shadows is imperative for any successful photometric stereo method. To estimate cast shadows for a ray-surface intersection point $\boldsymbol{\mathbf{X}}_{I}$, we raytrace from that point to the light source following the direction of the lighting vectors $\boldsymbol{\mathbf{L}}_{m}$ computed above. For each ray we take 16 random samples $j$ $\in[2\text{mm},50\text{mm}]$. For all these points, we query the SDF network and accumulate opacity following the same volumetric rendering procedure, i.e $s=\prod_{j}(1-\alpha_{j})$. We note that this shadow computation procedure is very computationally expensive therefore it cannot be computed for all the rendered points. Instead, we only compute it for the intersection point of each ray ($\boldsymbol{\mathbf{X}}_{I}$ is Figure 3) and assume it is the same for all other ray samples.
Visualisation of the generated renderings of synthetic and real data are shown in Figure 4.